49. If two fair dice are rolled, find the probability that the sum of the faces is 12.
Answer #1

Solution-

When two fair dice are rolled then

Number of total possible outcomes

= 6×6 = 36

Number of favorable outcomes

=Number of times the sum of numbers of the faces dice is 12

= 1

(Only one outcomes i.e (6,6) )

So,

The Probability that sum on the faces is 12

= Number of favorable outcomes/total number of possible outcomes

=1/36

Hence, Probability that sum on the faces of two fair rolled dice to be 12 is 1/36 .